# Seeing Eye to AI: Human Alignment via Gaze-Based Response Rewards for Large Language Models

### v1

@@ 4.2 Results
5 Related Work
Reward Modelling. The most popular approach to reward modeling follows the framework introduced by Ouyang et al. (2022). Several studies have examined alternative versions for refining RM s. For instance, Bai et al. (2022a) proposed more fine-grained reward structures, evaluating helpfulness and harmlessness separately. Other approaches have explored different reward modelling strategies (Wu et al., 2023a; Dong et al., 2023a; Wang et al., 2023c). Another line of research has focused on Process Based Reward Models (PRMs)(Lightman et al., 2024; Uesato et al., 2022) which differ from conventional RM s by predicting the correctness of intermediate steps, rather than solely evaluating final outputs. Other studies implement data augmentation techniques (Shen et al., 2023), or cross-attention mechanisms between encoded input text and candidate pairs (Jiang et al., 2023b). Moreover, some works have leveraged synthetic preference data for reward modelling (Cui et al., 2024; Jiao, 2023). Wu et al. (2024a) built upon the LLM-as-a-Judge framework Zheng et al. (2023) by introducing LLM-as-a-Meta-Judge, which evaluates the model’s judgments to generate preference pairs that enhance its decision-making capabilities. Finally, Pace et al. (2024) incorporated a self-training approach to improve reward model training. However, to date, no research has explored the integration of ET or other implicit feedback signals into RM.
Eye-tracking in Natural Language Processing. Several studies have investigated the use of ET data for a variety of NLP tasks, such as named entity recognition (Hollenstein and Zhang, 2019; Ren and Xiong, 2023; Yu et al., 2024; Hollenstein et al., 2019), text comprehension (Ahn et al., 2020; Reich et al., 2022; Sood et al., 2020b), language modelling (Huang et al., 2023; Huang and Hollenstein, 2023; Deng et al., 2023b), and question answering (Zhang and Hollenstein, 2024; Wang et al., 2024a). Other applications include code comprehension (Alakmeh et al., 2024), code summarization (Zhang et al., 2024) and hallucination detection (Maharaj et al., 2023). Eye-tracking has also been applied to sentiment analysis and sarcasm detection tasks (Mishra et al., 2016, 2017, 2018; Barrett et al., 2018; Huang et al., 2023; Khurana et al., 2023; Hollenstein et al., 2019; Yang and Hollenstein, 2023; Kiegeland et al., 2024; Deng et al., 2023a; Mathias et al., 2018; McGuire and Tomuro, 2021). The most relevant work to our approach is by
Kiegeland et al. (2024), which introduced a dataset generation method using ET signals for DPO, building on the controlled sentiment generation framework proposed by Deng et al. (2023a); Yang and Hollenstein (2023). While this study has contributed to the first steps towards integrating ET for human alignment in LLMs, it is task- and dataset-specific, often relying on ranking criteria that underutilize the potential of ET feedback. In contrast, our approach presents a more general framework by directly incorporating implicit feedback into the RM, rather than limiting its application to dataset creation.
@@ 6.1 Limitations & Future Work
Data. A limitation of our study is that both ET prediction models were trained on a relatively small datasets (Appendix A.1.2) that are not tailored to our tasks. Future work could benefit from directly collecting ET data specifically for LLM-generated responses, to offer insights into human reading comprehension and information processing of prompts, which could further improve model performance. Additionally, since the ET prediction models used in our experiments were trained on English corpora, the method’s generalizability to other languages requires further investigation. Moreover, we explored two methodologies for integrating ET features into the RM, but other approaches could prove more effective. For instance, ET features could be used to modify the RM ’s attention mask, as suggested by Zhang and Hollenstein (2024). Regarding dataset selection, both models used, Mistral 7B (Jiang et al., 2023a) and Llama 3 (Dubey et al., 2024), were fine-tuned on publicly available data, though specific details on the datasets are limited. Therefore, we cannot discount the possibility that the datasets we used may have been part of the models’ pretraining, particularly for Llama 3 7B Instruct, which has already undergone human alignment. However, as we compare against baselines using the same model checkpoints, any potential effects would be consistent across both conditions.
@@
All the code necessary to reproduce this work will be released in a GitHub repository once published. Both datasets used are publicly available (Köpf et al., 2023; Wang et al., 2024b). Additionally, both ET prediction models have been trained with public datasets (Cop et al., 2017; Hollenstein et al., 2018, 2020b).
@@
Since our research uses only synthetic ET data, there are no privacy concerns or need for large-scale experiments involving human subjects.
@@ A.1 Implementation details
This section provides further details on the implementation of our method. Subsection A.1.1 provides more details on the datasets used and their preprocessing steps. In subsection A.1.2, further information is given about the models used for generating ET features, along with the specific preprocessing required for each. Subsection A.1.3 explains the process of mapping the fixations from the tokenizer used by the generation model to the tokenizer used by the Reward Model. Finally, additional implementation details are discussed in subsection A.1.4.
@@ A.1.2 Eye-tracker features generation models
First model: Model presented in Huang and Hollenstein [2023]. The code for the model along with the weights is publicly available, so we used the pre-trained checkpoint and we adapted their code for our implementation. This model was trained on several eye-tracking datasets, including Dundee [Kennedy et al., 2012], GECO [Cop et al., 2017], ZuCo1 [Hollenstein et al., 2018], ZuCo2 [Hollenstein et al., 2020b]. More detailed information about this datasets is presented in Table 6. In Figure 3 we show an example of the synthetic total reading time (TRT)generated for the chosen and rejected response to a prompt.
@@
Second model: Model presented in Li and Rudzicz [2021]. The code and training data are also publicly available, so we trained it following their original methodology and we adapted their code to incorporate it into our implementation. The model was trained using the ZuCo 1 [Hollenstein et al., 2018] and ZuCo 2 [Hollenstein et al., 2020b] and PROVO [Luke and Christianson, 2018] datasets. For the ZuCo datasets, 800 sentences (15.7 tokens) were provided as training data and 191 sentences (3.5k tokens) were held out for evaluation. Information about the datasets used to train these models is in Table 6.
@@ A.1.3 Mapping ET features between different tokenizers
For each predictor, we reverse the method used to convert word-level features into token-level features (more details in Appendix A.1.2) but passing from tokens in the first one, to tokens in the second tokenizer. For example, if for the first ET features predictor models tokens $t_{1},t_{2}$ are mapped to tokens $t_{1},t_{2},t_{3}$ in another second tokenizer, the values sum for all the tokens in the first list and distribute them equally across all the tokens in the second list: being $t_{1}$ (1s TRT) and $t_{2}$ (2s TRT) each of $t_{1},t_{2},t_{3}$ are assigned a TRT of $(1+2)/3=1s$. In Table 7 is represented a example of this process where row TRT(1) are the finatl TRT mapped for the fist ET predictor and TRT(2) for the second one. Finally, because special chat tokens were removed when generating the ET features, we assign value 00 for all features in this tokens. At the time of publishing this work, some of the tokenizer functionalities needed for alignment between tokenizers were not available in Mistral 7B.
@@
A.1.4 Training datails
During training, the ET features predictor model remains frozen, while the ET features projector is normally trained without adapters (Figure 2). The RM is fine-tuned on top of the open-source models using QLoRA [Dettmers et al., 2023] based on Low-Rank Adaptation (LoRA)[Hu et al., 2021], which fine-tunes select dense layers by optimizing low-rank decomposition matrices representing weight changes, instead of directly updating pre-trained weights. QLoRA introduces memory optimization techniques such as the 4-bit NormalFloat (NF4), a novel data type, to improve performance without increasing memory usage. Following Dettmers et al. [2023] we use hyperparameters: r=8, alpha=32, and dropout=0.1.
@@ Appendix C Hyper parameter tuning
ET features projector The PyTorch architecture of our ET features projector is shown below. $num\_features$ varies between 1, 5, and 2 (depending on the configuration used $fcomb_{1}$, $fcomb_{2.5}$, $fcomb_{2.2}$ subsection 4.2). $p_{1},p_{2}$ are dropout values. Finally, after testing different combinations, we used $0.1$ and $0.3$.
@@
(a) Validation Loss Comparison for different LR Schedulers
(b) Learning rate comparison for different LR Schedulers
(a) Validation Loss Comparison for different LR Schedulers
(b) Learning rate comparison for different LR Schedulers

### v2

@@ 4.2 Results
5 Discussion
In this work, we introduced a novel framework for integrating implicit feedback into the Reward Model, a key component for aligning LLMs and generating synthetic data for further alignment. We validated our approach using widely-adopted, open-source models such as Llama 3 and Mistral, for initializing the RM. By employing two different models to generate ET features, our results show that incorporating implicit feedback consistently improves the RM’s ability to predict user preferences, regardless of the model used and without the need to reach large parameter counts or train on massive datasets. Additionally, our method leverages ET features generated by models, making it fully scalable and applicable to various human alignment methods, including those that involve artificially generated datasets. This work advances the ongoing discussion on optimizing AI alignment with human values and shows the potential of multimodal signals for NLP research enhancing current methods.
5.1 Limitations & Future Work
Data. A limitation of our study is that both ET prediction models were trained on a relatively small datasets (Appendix A.1.2) that are not tailored to our tasks. Future work could benefit from directly collecting ET data specifically for LLM-generated responses, to offer insights into human reading comprehension and information processing of prompts, which could further improve model performance. Additionally, since the ET prediction models used in our experiments were trained on English corpora, the method’s generalizability to other languages requires further investigation. Moreover, we explored two methodologies for integrating ET features into the RM, but other approaches could prove more effective. For instance, ET features could be used to modify the RM ’s attention mask, as suggested by Zhang and Hollenstein (2024). Regarding dataset selection, both models used, Mistral 7B and Llama 3, were fine-tuned on publicly available data, though specific details on the datasets are limited. Therefore, we cannot discount the possibility that the datasets we used may have been part of the models’ pretraining, particularly for Llama 3 7B Instruct, which has already undergone human alignment. However, as we compare against baselines using the same model checkpoints, any potential effects would be consistent across both conditions. Finally, it is important to note that although the ET prediction models remain fixed during training, our solution has a slightly higher number of parameters than the baseline.
Training. The scaling trends for the RM (Touvron et al., 2023) show that larger models or models trained on massive datasets perform better. A promising direction would be to validate our framework on larger models, without relying on PEFT methods, and on larger datasets. We are confident that, despite the considerable computational costs this may entail, our framework is capable of scaling effectively. Another direction is integrating the proposed RM into an alignment method like RLHF, or applying it in rejection sampling to generate synthetic preference datasets, ensuring that accuracy gains in the RM translate to improvements in the final LLMs.
@@ 6.1 Limitations & Future Work
Data. A limitation of our study is that both ET prediction models were trained on a relatively small datasets (Appendix A.1.2) that are not tailored to our tasks. Future work could benefit from directly collecting ET data specifically for LLM-generated responses, to offer insights into human reading comprehension and information processing of prompts, which could further improve model performance. Additionally, since the ET prediction models used in our experiments were trained on English corpora, the method’s generalizability to other languages requires further investigation. Moreover, we explored two methodologies for integrating ET features into the RM, but other approaches could prove more effective. For instance, ET features could be used to modify the RM ’s attention mask, as suggested by Zhang and Hollenstein (2024). Regarding dataset selection, both models used, Mistral 7B (Jiang et al., 2023) and Llama 3 (Dubey et al., 2024), were fine-tuned on publicly available data, though specific details on the datasets are limited. Therefore, we cannot discount the possibility that the datasets we used may have been part of the models’ pretraining, particularly for Llama 3 7B Instruct, which has already undergone human alignment. However, as we compare against baselines using the same model checkpoints, any potential effects would be consistent across both conditions.
@@
All the code necessary to reproduce is in the GitHub repository  [^3] https://github.com/Telefonica-Scientific-Research/gaze_reward. Both datasets used are publicly available (Köpf et al., 2023; Wang et al., 2024b). Additionally, both ET prediction models have been trained with public datasets (Cop et al., 2017; Hollenstein et al., 2018, 2020b).
@@
Since our research uses only synthetic ET data, there are no privacy concerns or need for large-scale experiments involving human subjects. We should also raise attention to the limitations of human feedback and ET prediction models bias, that inevitably reflect aspects of their training data.
Acknowledgments
This research is supported by Horizon Europe’s European Innovation Council through the Pathfinder program (SYMBIOTIK project, grant 101071147) and by the Industrial Doctorate Plan of the Department of Research and Universities of the Generalitat de Catalunya, under Grant AGAUR 2023 DI060. We also want to thank Sebastian Macaluso for his important feedback during the project.
@@ A.1 Implementation details
This section provides further details on the implementation of our method. Subsection A.1.1 provides more details on the datasets used and their preprocessing steps. In subsection A.1.2, further information is given about the models used for generating ET features, along with the specific preprocessing required for each. Subsection A.1.3 explains the process of mapping the fixations from the tokenizer used by the generation model to the tokenizer used by the Reward Model. Subsection A.1.4 give more detials on the checkpoints used fot the RM backbone models. Finally, additional implementation details are discussed in subsection A.1.5.
@@ A.1.2 Eye-tracker features generation models
First model: Model presented in Huang and Hollenstein [2023]. The code for the model along with the weights is publicly available, so we used the pre-trained checkpoint and we adapted their code for our implementation. This model was trained on several eye-tracking datasets, including Dundee [Kennedy et al., 2012], GECO [Cop et al., 2017], ZuCo1 [Hollenstein et al., 2018], ZuCo2 [Hollenstein et al., 2020b]. More detailed information about this datasets is presented in Table 7. The best model achieves an mean squared error (MSE)of 4.02 on a randomly held-out test set (25% of all data). This model has 17.5M pararemetes and remains frozen during training. In Figure 3 we show an example of the synthetic total reading time (TRT)generated for the chosen and rejected response to a prompt.
@@
Second model: Model presented in Li and Rudzicz [2021]. The code and training data are also publicly available, so we trained it following their original methodology and we adapted their code to incorporate it into our implementation. The model was trained using the ZuCo 1 [Hollenstein et al., 2018] and ZuCo 2 [Hollenstein et al., 2020b] and PROVO [Luke and Christianson, 2018] datasets. For the ZuCo datasets, 800 sentences (15.7 tokens) were provided as training data and 191 sentences (3.5k tokens) were held out for evaluation. Information about the datasets used to train these models is in Table 7. The model is based on RoBERTa [Liu et al., 2019] with a regression head on each token. This head is a linear layer that outputs five features: FFD, fixProp, GPT, TRT, and nFix (Table 1). mean absolute error (MAE)for each feature is presented in Table 6. This model has 125M parameters and remains frozen during training.
@@
Table 6: MAE performance of the model reported in Li and Rudzicz [2021].
nFix FFD GPT TRT fixProp All (Dev)
3.984 0.713 2.424 1.556 10.781 3.892
@@ A.1.3 Mapping ET features between different tokenizers
For each predictor, we reverse the method used to convert word-level features into token-level features (more details in Appendix A.1.2) but passing from tokens in the first one, to tokens in the second tokenizer. For example, if for the first ET features predictor models tokens $t_{1},t_{2}$ are mapped to tokens $t_{1},t_{2},t_{3}$ in another second tokenizer, the values sum for all the tokens in the first list and distribute them equally across all the tokens in the second list: being $t_{1}$ (1s TRT) and $t_{2}$ (2s TRT) each of $t_{1},t_{2},t_{3}$ are assigned a TRT of $(1+2)/3=1s$. In Table 8 is represented a example of this process where row TRT(1) are the final TRT mapped for the first ET predictor and TRT(2) for the second one. Finally, because special chat tokens were removed when generating the ET features, we assign value 00 for all features in this tokens. At the time of publishing this work, some of the tokenizer functionalities needed for alignment between tokenizers were not available in Mistral 7B.
@@
A.1.4 Models
As RM base models we use the pretrained checkpoint of Hugging Face for Llama 3 8B  [^9] https://huggingface.co/meta-llama/Meta-Llama-3-8B [Dubey et al., 2024], Llama 3 8B-instruct  [^10] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct[Dubey et al., 2024] and Mistral 7B  [^11] https://huggingface.co/mistralai/Mistral-7B-v0.3 [Jiang et al., 2023].
A.1.5 Training datails
During training, the ET features predictor model remains frozen. The RM is fine-tuned on top of the open-source models using QLoRA [Dettmers et al., 2023] based on Low-Rank Adaptation (LoRA)[Hu et al., 2021], which fine-tunes select dense layers by optimizing low-rank decomposition matrices representing weight changes, instead of directly updating pre-trained weights. QLoRA introduces memory optimization techniques such as the 4-bit NormalFloat (NF4), a novel data type, to improve performance without increasing memory usage. Following Dettmers et al. [2023] we use hyperparameters: r=8, alpha=32, and dropout=0.1.
@@ Appendix C Hyper parameter tuning
(a) Validation Loss Comparison for different LR Schedulers
(b) Learning rate comparison for different LR Schedulers
(a) Validation Loss Comparison for different LR Schedulers
(b) Learning rate comparison for different LR Schedulers
ET features projector The PyTorch architecture of our ET features projector is shown below. $num\_features$ varies between 1, 5, and 2 (depending on the configuration used $fcomb_{1}$, $fcomb_{2.5}$, $fcomb_{2.2}$ subsection 4.2). $p_{1},p_{2}$ are dropout values. Finally, after testing different combinations, we used $0.1$ and $0.3$. This model has 0.53M parameters.